# On the effects of radiation on mass transfer in binary stars

### v1

@@ Abstract
Mass transfer (MT) in binary systems is a common evolutionary process that can significantly affect the structure, evolution, and final fate of both stars. In modeling MT hydrodynamics, it is usually assumed that the critical point of the flow, where the velocity exceeds the local sound speed, coincides with the inner Lagrange point (L1). However, in massive donors, radiative pressure dominates over gas pressure and the Eddington factor $\Gamma_{\text{Edd}}$ can reach or exceed unity. If there is significant momentum and/or energy exchange between gas and radiation, the critical point of the flow can shift away from L1, affecting the MT rate ($\dot{M}_{\text{d}}$). Here, we investigate the effects of radiation on MT using time-steady radiative hydrodynamic equations and the assumption of the von Zeipel theorem. We provide analytical expressions that closely approximate $\dot{M}_{\text{d}}$, algebraic solutions for several simplified situations, and numerical solutions using a realistic equation of state. We find two main differences with respect to the traditional expressions for $\dot{M}_{\text{d}}$. First, for Roche-lobe-underfilling donors with $\Gamma_{\text{Edd}}\lesssim 1$, radiative momentum exchange leads to an exponential increase of $\dot{M}_{\text{d}}$ as a function of $1-\Gamma_{\text{Edd}}$. We provide a simple modification of existing prescriptions that takes this effect into account. Second, we find that the photon tiring limit for super-Eddington outflows is much less restrictive near L1 than in spherical stars. We suggest that donors with super-Eddington convectively inefficient subsurface layers are able to drive MT with $-\dot{M}_{\text{d}}\gtrsim 10^{-2}\,\rm{M}_{\odot}\,\text{yr}^{-1}$ even before overfilling their Roche lobes. We characterize the conditions for this new mode of super-Eddington-boosted MT and discuss the implications for binary evolution including the potential link to nonterminal outbursts of Luminous Blue Variables.
@@
pubyear: 2025pagerange: On the effects of radiation on mass transfer in binary starsE
@@ 1 Introduction
Massive stars drive cosmic evolution through their winds, photons, and explosions. Their evolutionary journey ultimately leads to the synthesis of new chemical elements and the formation of neutron stars and black holes. Most massive stars exist in gravitationally bound binaries or higher-order multiples (e.g. Moe & Di Stefano, 2017) and mutual interaction between stars can significantly alter their evolution, ultimate fate, and imprint on surroundings (e.g. Sana et al., 2012; Tauris & van den Heuvel, 2023; Marchant & Bodensteiner, 2024). Binary stars often interact through mass transfer (MT), where material flows from one star through the first Lagrange point (L1) to its companion. Ongoing MT can be detected by observing the energy released as the material falls toward the companion or by identifying signatures of diffuse gas surrounding the binary system when some material escapes the binary completely. Over time, MT can significantly alter both the internal structures and external appearances of stars, making evolutionary paths impossible for isolated stars. In some cases, MT becomes unstable, triggering common envelope evolution, which is a critical phase in many evolutionary pathways that lead to compact object binaries and gravitational wave progenitors (Paczynski, 1976; Belczynski et al., 2002; Mandel & Broekgaarden, 2022; Röpke & De Marco, 2023).
@@
One key difference between massive stars and low-mass stars is the significant role of radiation pressure. The increasing importance of radiation pressure with stellar mass, both inside and outside the photosphere, manifests in relations between global properties of stars, such as the mass–luminosity relation on the main sequence, the intensification of line-driven winds, and positively contributes to the viability of phenomena such as chemically homogeneous evolution (e.g., Maeder & Meynet, 1987; Maeder, 1987; Heger & Langer, 2000; Vink et al., 2000; Yoon et al., 2006).
@@
Massive stars can exceed the continuum Eddington limit, where the outward frequency-integrated radiative acceleration exceeds the inward gravitational pull. The existence and position of super-Eddington regions depend on density, temperature, and opacity profiles inside the star. Historically, approaching or exceeding the Eddington limit has been argued to be associated with increased mass loss. There is observational evidence that the proximity to the Eddington limit is the physical reason for the onset of Wolf-Rayet-type mass loss (e.g. Gräfener et al., 2011) and a potential mechanism responsible for luminous blue variable (LBV) eruptions (e.g., Lamers & Fitzpatrick, 1988; Smith & Owocki, 2006; Smith et al., 2011).
@@
The simple fact of existence of a super-Eddington region is not sufficient to generate outflows because the energy flux can instead be carried out by convection (Joss et al., 1973). Near the stellar surface, convection can become inefficient once the required convective velocities approach or exceed the local sound speed. One-dimensional stellar models are able to support limited convectively inefficient super-Eddington regions by means of envelope inflation, density and pressure inversions, and possibly pulsations (e.g., Glatzel & Kiriakidis, 1993; Sanyal et al., 2015). Owocki et al. (2004) argued that the mass-loss rate of any super-Eddington wind driven from the region of inefficient convection would far exceed the photon tiring limit, where all available stellar luminosity is used to lift material from the star to infinity, and such outflow would stagnate at some finite radius and fall back to the star. Consequently, multidimensional effects such as hydrodynamic instabilities and effectively porous envelopes for radiation transfer are likely to significantly alter these spherically symmetric predictions (e.g., Mihalas, 1969; Shaviv, 2001b; Begelman, 2001; Owocki et al., 2004). This was confirmed by time-dependent simulations of Jiang et al. (2015, 2018), which showed complex outflow and inflow patterns with occasional eruptions. Nevertheless, continuum super-Eddington winds can still occur if some additional energy is deposited below the surface, and the photon-tiring limit remains important in these situations (Shaviv, 2001a; Owocki et al., 2004, 2017; Quataert et al., 2016; Shen et al., 2016).
@@
The ideas and results on super-Eddington regions in stars rely on the crucial fact that radiative and gravitational accelerations in a spherically symmetric object have the same inverse-square scaling with radius. This is not the case for Roche-lobe-filling stars in binaries, where effective gravity vanishes at the L1 point. Consequently, in a fraction of the stellar envelope, radiative forces can potentially overwhelm gravity and drive much higher MT rates because the photon tiring now needs to overcome much smaller gravitational potential difference. Such a super-Eddington MT boost could play an important role in binary evolution involving massive stars and for explaining the origin of eruptions in LBVs, which are or could have once been a component of a binary system (e.g., Smith & Owocki, 2006; Smith & Tombleson, 2015; Mahy et al., 2022).
@@
In this paper, our objective is to characterize the effects of radiation pressure on MT through the L1 point. In particular, we are interested in understanding MT from donors with a super-Eddington layer located immediately inside the L1 point, as illustrated in Fig. 1. Our approach builds on our previous work in Cehula & Pejcha (2023, hereafter CP23), where we averaged the three-dimensional equations of hydrodynamics in the plane normal to the axis connecting both stars and solved the resulting one-dimensional two-point boundary value problem. In other words, we solved the structure of the star just interior the L1 point allowing for velocity terms and connecting to the hydrostatic spherically symmetric stellar model at the inner boundary. Here, we modify the method to include radiation. By finding a steady hydrodynamic solution in a gravitational potential near L1, we are in principle able to isolate effects that are inaccessible in the traditional approach, where the MT rate is calculated by post-processing a spherically symmetric stellar model (e.g. Ritter, 1988; Kolb & Ritter, 1990; Jackson et al., 2017; Marchant et al., 2021).
@@ 2.1 General three-dimensional equations
We start with stationary frequency-integrated 3D radiation hydrodynamic equations assuming that radiation has a blackbody spectrum at the local radiation temperature, scattering is negligible, flux mean opacity is equal to Rosseland mean opacity (which is valid in optically thick regions), and using flux-limited diffusion approximation (e.g. Alme & Wilson, 1973; Mihalas & Mihalas, 1984; Krumholz et al., 2007). The mixed-frame formulation correct to $\mathcal{O}(v/c)$ in the streaming, static diffusion, and dynamic diffusion regimes including the Roche potential $\phi_{\rm R}$ reads (e.g. Zhang et al., 2011; Calderón et al., 2021)
@@
where $\epsilon_{\rm gas}$ is the gas (internal) specific energy in units of $\text{erg}\>\text{g}^{-1}$. The flux limiter $\lambda$ and the Eddington factor $f$ are defined by (Levermore & Pomraning, 1981)
@@
Radiation energy density in the lab frame and the co-moving frame are related by (Zhang et al., 2011)
@@
Based on these assumptions, we can derive a velocity equation for ideal gas in the form following Lubow & Shu (1975, eq. 10), which elucidates the terms contributing to the existence and position of the critical point,
@@
and where we subtract the advective flux $\boldsymbol{F}_{\rm adv}\equiv\boldsymbol{v}E_{\rm rad}+\boldsymbol{v\cdot}\boldsymbol{\mathsf{P}}_{\rm rad}$ from the lab frame radiative flux $\boldsymbol{F}_{\rm rad}$, where $\boldsymbol{\mathsf{P}}_{\rm rad}$ is the radiation pressure tensor (Krumholz et al., 2007). The gas adiabatic sound speed $c_{s,\rm gas}$ reads
@@
There are three differences between our velocity equation (6) and the isothermal velocity equation of Lubow & Shu (1975, their eq. 10). First, our velocity equation includes the term $\boldsymbol{v\cdot}\boldsymbol{f}_{\rm rad}$ that accounts for the increase in radiation momentum originating from the radiation pressure. Second, our equation involves the term $-\left(\gamma-1\right)\left|\boldsymbol{v}\right|q_{\rm rad}$ that accounts for the gas specific heating/cooling rate (in units of $\text{erg}\>\text{g}^{-1}\>\text{s}^{-1}$) due to energy exchange between the gas and radiation. Third, our critical velocity is the gas adiabatic sound speed $c_{s,\rm gas}$ instead of the gas isothermal sound speed $c_{T,\rm gas}$.
@@ 2.2 General one-dimensional equations
Let us now assume that the gas flows through a nozzle oriented along the $x$ axis and with a cross-section denoted by $Q(x)$. In CP23, we calculated $Q(x)$ based on the Roche potential and thermodynamic properties of the gas flowing to the L1 point. Assuming that all quantities are constant at a given $Q(x)$, we can reduce the set of 3D equations (1) to the set of 1D equations (see Appendix B for details),
@@
and equation (1f) remains unchanged. The donor’s mass loss rate is given by (see also CP23, )
$-\dot{M}_{\rm d}=v\rho Q.$ (13)
@@
which is the general form of the momentum equation of a spherically symmetric stellar wind with energy input $\left(q_{\rm rad}\right)$ and momentum input $\left(f_{\rm rad}\right)$ (Lamers & Cassinelli, 1999, eq. 4.18). Moreover, considering equation (11d), we get $f_{\rm rad}=(L_{\rm d,rad}/L_{\rm d,Edd})(GM_{\rm d}/r^{2})=(L_{\rm d,rad}/L_{\rm d,Edd})({\rm d}\phi_{\rm R}/{\rm d}r)$, where $L_{\rm d,rad}$ is the donor’s radiative (diffusive) luminosity and $L_{\rm d,Edd}$ is the donor’s Eddington luminosity and we have used $L_{\rm N}=L_{\rm d}$, where $L_{\rm d}$ is the donor’s total luminosity and $L_{\rm d,Edd}=4\pi cGM_{\rm d}/\kappa_{\rm R}$.
@@ 2.3 The von Zeipel theorem
The set of 1D radiation hydrodynamic equations (11) can be further simplified. We can use an extension of the von Zeipel theorem, which states that the radiative flux in a uniformly rotating star is proportional to the local effective gravity (von Zeipel, 1924), and write for the donor’s envelope (see also recent works by Fabry et al., 2022, 2023)
@@
The fundamental assumption underlying von Zeipel approximation is that the donor is barotropic, i.e. $\rho$, $T$, etc. are only functions of $\phi_{\rm R}$ (von Zeipel, 1924) and that the effects of convection are ignored. The application of this theorem to Roche-lobe-filling stars also seems problematic, because the von Zeipel theorem predicts zero radiative flux at the L1 point. We will further discuss the limitations in Section 6.
@@ 3.1 Optically-thick hydrostatic case
Using the above set of equations, we can in principle reconstruct part of the donor’s hydrostatic envelope close to the L1 point, specifically, between some point $x_{0}$ and the L1 point located at $x_{1}$, $x_{0}<x_{1}$. In order to close the system of equations we still need an EOS for $P_{\rm gas}$ and two boundary conditions: $\rho(x_{0})=\rho_{0}$ and $T(x_{0})=T_{0}$ that need to be taken from the underlying hydrostatic structure of a stellar evolution code. We also need to evaluate the Eddington ratio $\Gamma_{\rm Edd}$ (equation 17). The exact way to connect the $x$ coordinate with the potential $\phi$ and consecutively with the radius $r$ was explained in CP23 and will also be discussed later in Section 4.1.
@@ 3.2 Optically-thick, $v\ll c$ case
We can generalize the previous case to a situation where there is a slow but nonzero velocity, $v\ll c$. We apply assumptions similar to those in the hydrostatic case , but in order to obtain the continuity equation, we need a prescription for the cross section of the nozzle $Q$. In our previous work (CP23, but see also Meyer & Meyer-Hofmeister 1983), we assumed $Q\propto c_{T,\rm gas}^{2}\propto P_{\rm gas}/\rho$, which assumes that the gas is in adiabatic hydrostatic equilibrium in the plane perpendicular to to flow toward L1. For simplicity, we apply the same assumption here and we arrive at the following set of equations

### v2

@@ Abstract
Mass transfer (MT) in binary systems is a common evolutionary process that can significantly affect the structure, evolution, and final fate of both stars. In modeling MT hydrodynamics, it is usually assumed that the critical point of the flow, where the velocity exceeds the local sound speed, coincides with the inner Lagrange point (L1). However, in massive donors where radiative pressure dominates over gas pressure and the Eddington factor $\Gamma_{\text{Edd}}$ can approach or exceed unity, radiation–gas coupling can shift the critical point away from L1, altering the MT rate ($\dot{M}_{\text{d}}$). We investigate the effects of radiation on MT using time-steady radiative hydrodynamic equations and the von Zeipel theorem. We derive analytical expressions that closely approximate $\dot{M}_{\text{d}}$, algebraic solutions for simplified cases, and numerical results using a realistic equation of state. Two main differences emerge relative to traditional prescriptions for $\dot{M}_{\text{d}}$. First, for Roche-lobe-underfilling donors with $\Gamma_{\text{Edd}}\lesssim 1$, radiative momentum exchange leads to an exponential increase of $\dot{M}_{\text{d}}$ as a function of $1-\Gamma_{\text{Edd}}$. We provide a simple modification of existing prescriptions that captures this effect. Second, the photon tiring limit for super-Eddington outflows is much less restrictive near L1 than in spherical stars. We suggest that donors with super-Eddington, convectively inefficient subsurface layers can drive MT with $\dot{M}_{\text{d}}\gtrsim 10^{-2}\,\rm{M}_{\odot}\,\text{yr}^{-1}$ even before Roche-lobe overflow. We characterize the conditions for this new mode of super-Eddington-boosted MT and discuss its implications for binary evolution, including potential links to nonterminal outbursts of Luminous Blue Variables.
@@
pubyear: 2026pagerange: On the effects of radiation on mass transfer in binary starsE
@@ 1 Introduction
Massive stars drive cosmic evolution through their winds, photons, and explosions. Their evolutionary journey ultimately leads to the synthesis of new chemical elements and the formation of neutron stars and black holes. Most massive stars exist in gravitationally bound binaries or higher-order multiples (e.g. moe2017) and mutual interaction between stars can significantly alter their evolution, ultimate fate, and imprint on surroundings (e.g. sana2012; tauris2023; marchant2024). Binary stars often interact through mass transfer (MT), where material flows from one star through the first Lagrange point (L1) to its companion. Ongoing MT can be detected by observing the energy released as the material falls toward the companion or by identifying signatures of diffuse gas surrounding the binary system when some material escapes the binary completely. Over time, MT can significantly alter both the internal structures and external appearances of stars, making evolutionary paths impossible for isolated stars. In some cases, MT becomes unstable, triggering common envelope evolution, which is a critical phase in many evolutionary pathways that lead to compact object binaries and gravitational wave progenitors (paczynski1976; belczynski2002; mandel2022; roekpke2023).
@@
One key difference between massive stars and low-mass stars is the significant role of radiation pressure. The increasing importance of radiation pressure with stellar mass, both inside and outside the photosphere, manifests in relations between global properties of stars, such as the mass–luminosity relation on the main sequence, the intensification of line-driven winds, and positively contributes to the viability of phenomena such as chemically homogeneous evolution (e.g., maeder_meynet87; maeder87; heger00; vink00; yoon06).
@@
Massive stars can exceed the continuum Eddington limit, where the outward frequency-integrated radiative acceleration exceeds the inward gravitational pull. The existence and position of super-Eddington regions depend on density, temperature, and opacity profiles inside the star. Historically, approaching or exceeding the Eddington limit has been argued to be associated with increased mass loss. There is observational evidence that the proximity to the Eddington limit is the physical reason for the onset of Wolf-Rayet-type mass loss (e.g. grafener2011) and a potential mechanism responsible for luminous blue variable (LBV) eruptions (e.g., lamers88; smith_owocki06; smith2011).
@@
The simple fact of existence of a super-Eddington region is not sufficient to generate outflows because the energy flux can instead be carried out by convection (joss73). Near the stellar surface, convection can become inefficient once the required convective velocities approach or exceed the local sound speed. One-dimensional stellar models are able to support limited convectively inefficient super-Eddington regions by means of envelope inflation, density and pressure inversions, and possibly pulsations (e.g., glatzel93; sanyal2015). owocki2004 argued that the mass-loss rate of any super-Eddington wind driven from the region of inefficient convection would far exceed the photon tiring limit, where all available stellar luminosity is used to lift material from the star to infinity, and such outflow would stagnate at some finite radius and fall back to the star. Consequently, multidimensional effects such as hydrodynamic instabilities and effectively porous envelopes for radiation transfer are likely to significantly alter these spherically symmetric predictions (e.g., mihalas69; shaviv01a; begelman01; owocki2004). This was confirmed by time-dependent simulations of jiang2015; jiang2018, which showed complex outflow and inflow patterns with occasional eruptions. Nevertheless, continuum super-Eddington winds can still occur if some additional energy is deposited below the surface, and the photon-tiring limit remains important in these situations (shaviv01b; owocki2004; owocki2017; quataert16; shen16).
@@
The ideas and results on super-Eddington regions in stars rely on the crucial fact that radiative and gravitational accelerations in a spherically symmetric object have the same inverse-square scaling with radius. This is not the case for Roche-lobe-filling stars in binaries, where effective gravity vanishes at the L1 point. Consequently, in a fraction of the stellar envelope, radiative forces can potentially overwhelm gravity and drive much higher MT rates because the photon tiring now needs to overcome much smaller gravitational potential difference. Such a super-Eddington MT boost could play an important role in binary evolution involving massive stars and for explaining the origin of eruptions in LBVs, which are or could have once been a component of a binary system (e.g., smith_owocki06; smith2015; mahy2022).
@@
In this paper, our objective is to characterize the effects of radiation pressure on MT through the L1 point. In particular, we are interested in understanding MT from donors with a super-Eddington layer located immediately inside the L1 point, as illustrated in Fig. 1. Our approach builds on our previous work in cehula2023, where we averaged the three-dimensional equations of hydrodynamics in the plane normal to the axis connecting both stars and solved the resulting one-dimensional two-point boundary value problem. In other words, we solved the structure of the star just interior the L1 point allowing for velocity terms and connecting to the hydrostatic spherically symmetric stellar model at the inner boundary. Here, we modify the method to include radiation. By finding a steady hydrodynamic solution in a gravitational potential near L1, we are in principle able to isolate effects that are inaccessible in the traditional approach, where the MT rate is calculated by post-processing a spherically symmetric stellar model (e.g. ritter1988; kolb1990; jackson2017; marchant2021).
@@ 2.1 General three-dimensional equations
We start with stationary frequency-integrated 3D radiation hydrodynamic equations assuming that radiation has a blackbody spectrum at the local radiation temperature, scattering is negligible, flux mean opacity is equal to Rosseland mean opacity (which is valid in optically thick regions), and using flux-limited diffusion approximation (e.g. alme1973; mihalas1984; krumholz2007). The mixed-frame formulation correct to $\mathcal{O}(v/c)$ in the streaming, static diffusion, and dynamic diffusion regimes including the Roche potential $\phi_{\rm R}$ reads (e.g. zhang2011; calderon2021)
@@
where $\epsilon_{\rm gas}$ is the gas (internal) specific energy in units of $\text{erg}\>\text{g}^{-1}$. The flux limiter $\lambda$ and the Eddington factor $f$ are defined by (levermore1981)
@@
Radiation energy density in the lab frame and the co-moving frame are related by (zhang2011)
@@
Based on these assumptions, we can derive a velocity equation for ideal gas in the form following lubow1975, which elucidates the terms contributing to the existence and position of the critical point,
@@
and where we subtract the advective flux $\boldsymbol{F}_{\rm adv}\equiv\boldsymbol{v}E_{\rm rad}+\boldsymbol{v\cdot}\boldsymbol{\mathsf{P}}_{\rm rad}$ from the lab frame radiative flux $\boldsymbol{F}_{\rm rad}$, where $\boldsymbol{\mathsf{P}}_{\rm rad}$ is the radiation pressure tensor (krumholz2007). The gas adiabatic sound speed $c_{s,\rm gas}$ reads
@@
There are three differences between our velocity equation (6) and the isothermal velocity equation of lubow1975. First, our velocity equation includes the term $\boldsymbol{v\cdot}\boldsymbol{f}_{\rm rad}$ that accounts for the increase in radiation momentum originating from the radiation pressure. Second, our equation involves the term $\left(\gamma-1\right)\left|\boldsymbol{v}\right|q_{\rm rad}$ that accounts for the gas specific heating/cooling rate (in units of $\text{erg}\>\text{g}^{-1}\>\text{s}^{-1}$) due to energy exchange between the gas and radiation. Third, our critical velocity is the gas adiabatic sound speed $c_{s,\rm gas}$ instead of the gas isothermal sound speed $c_{T,\rm gas}$.
@@ 2.2 General one-dimensional equations
Let us now assume that the gas flows through a nozzle oriented along the $x$ axis and with a cross-section denoted by $Q(x)$. In cehula2023, we calculated $Q(x)$ based on the Roche potential and thermodynamic properties of the gas flowing to the L1 point. Assuming that all quantities are constant at a given $Q(x)$, we can reduce the set of 3D equations (1) to the set of 1D equations (see Appendix B for details),
@@
and equation (1f) remains unchanged. The donor’s mass loss rate is given by (see also cehula2023, )
$\dot{M}_{\rm d}=v\rho Q.$ (13)
@@
which is the general form of the momentum equation of a spherically symmetric stellar wind with energy input $\left(q_{\rm rad}\right)$ and momentum input $\left(f_{\rm rad}\right)$ (lamers1999, eq. 4.18). Moreover, considering equation (11d), we get $f_{\rm rad}=(L_{\rm d,rad}/L_{\rm d,Edd})(GM_{\rm d}/r^{2})=(L_{\rm d,rad}/L_{\rm d,Edd})({\rm d}\phi_{\rm R}/{\rm d}r)$, where $L_{\rm d,rad}$ is the donor’s radiative (diffusive) luminosity and $L_{\rm d,Edd}$ is the donor’s Eddington luminosity and we have used $L_{\rm N}=L_{\rm d}$, where $L_{\rm d}$ is the donor’s total luminosity and $L_{\rm d,Edd}=4\pi cGM_{\rm d}/\kappa_{\rm R}$.
@@ 2.3 The von Zeipel theorem
The set of 1D radiation hydrodynamic equations (11) can be further simplified. We can use an extension of the von Zeipel theorem, which states that the radiative flux in a uniformly rotating star is proportional to the local effective gravity (vonZeipel1924), and write for the donor’s envelope (see also recent works by fabry2022; fabry2023)
@@
The fundamental assumption underlying von Zeipel approximation is that the donor is barotropic, i.e. $\rho$, $T$, etc. are only functions of $\phi_{\rm R}$ (vonZeipel1924) and that the effects of convection are ignored. The application of this theorem to Roche-lobe-filling stars also seems problematic, because the von Zeipel theorem predicts zero radiative flux at the L1 point. We will further discuss the limitations in Section 6.
@@ 3.1 Optically-thick hydrostatic case
Using the above set of equations, we can in principle reconstruct part of the donor’s hydrostatic envelope close to the L1 point, specifically, between some point $x_{0}$ and the L1 point located at $x_{1}$, $x_{0}<x_{1}$. In order to close the system of equations we still need an EOS for $P_{\rm gas}$ and two boundary conditions: $\rho(x_{0})=\rho_{0}$ and $T(x_{0})=T_{0}$ that need to be taken from the underlying hydrostatic structure of a stellar evolution code. We also need to evaluate the Eddington ratio $\Gamma_{\rm Edd}$ (equation 17). The exact way to connect the $x$ coordinate with the potential $\phi$ and consecutively with the radius $r$ was explained in cehula2023 and will also be discussed later in Section 4.1.
@@ 3.2 Optically-thick, $v\ll c$ case
We can generalize the previous case to a situation where there is a slow but nonzero velocity, $v\ll c$. We apply assumptions similar to those in the hydrostatic case , but in order to obtain the continuity equation, we need a prescription for the cross section of the nozzle $Q$. In our previous work (cehula2023, but see also meyer1983), we assumed $Q\propto c_{T,\rm gas}^{2}\propto P_{\rm gas}/\rho$, which assumes that the gas is in adiabatic hydrostatic equilibrium in the plane perpendicular to to flow toward L1. For simplicity, we apply the same assumption here and we arrive at the following set of equations